OpenStudy (anonymous):

Can anyone solve this question integral of e^(m sine inverse of x) this is an integration by parts question

OpenStudy (anonymous):

\[\int\limits e ^{m \sin^{x} }\]

OpenStudy (anonymous):

I would have difficulties expressing that in a textbox. You are missing "dx." So, you would use u = e^(glob) and dv = dx. Also, you can use the integrate function at wolframalpha.com. Just type integrate, then your function, and it will compute the desired indefinite integral.